At a price and quantity of 70 and 60 respectively ,is the monopolistic firm in a long run or short run equilibrium
1
Expert's answer
2021-07-26T11:42:03-0400

Assuming at a price of 70, Q=2800 and at a price of 60, Q=3000

% change in Q will be


"=\frac{(3000-2800)}{\frac{3000+2800}{2}}\times100"

"=\frac{200}{2900}\times100"


"=6.9"


% change in price

"=\frac{60-70}{\frac{60+70}{2}}×100"


"=\frac{-10}{65}×100"


"=-15.4"


Price elasticity of demand

"\frac{6.9}{15.4}"


"=0.45."


At 70 the firm is running in the short run while at 60 the firm is running in the long run